### v1

@@ 4 Methodology
In this section, we first introduce some basic concepts and methods of reservoir computing, as well as some theoretical analysis. In particular, we consider optimization by improving the orthogonality of the reservoir digraph matrix, which enhancing its memory capacity. To achieve this goal, the GLMY theory will be used. Subsequently, we further introduce the persistent GLMY homology of a digraph and the proposed approach for optimizing the reservoirs.
@@ 4.1 Reservoir computing and its theoretical analysis
Training the weight matrices in deep learning models is typically computationally expensive. In contrast to conventional neural networks, where the weights of all layers are trained, the reservoir computing (RC) architecture only requires the training of the readout layer’s weights. This characteristic significantly reduces the computational complexity, offering a computational efficiency advantage in certain applications. The general structure of the reservoir computing model is illustrated in Figure 3. It primarily consists of three components: the input layer, the reservoir, and the output layer. The input layer is responsible for encoding the incoming data and feeding it into the reservoir. The reservoir consists of a large, fixed, and recurrent network of neurons that project the input data into a high-dimensional dynamic space. The output layer, which is typically trained, maps the high-dimensional representations from the reservoir to the desired output. Generally, the input and output layers are fully connected layers, while the reservoir is represented by a weight matrix $W_{\text{r}}\in\mathbb{R}^{N\times N}$, where $N$ denotes the size of the reservoir. The weight matrix defines the connections between the neurons within the reservoir, determining the structure and dynamics of the recurrent network. Similarly, the input layer $L_{\text{in}}$ and the output layer $L_{\text{output}}$ are also described by their respective weight matrices. Specifically, the input layer is represented by the weight matrix $W_{\text{in}}\in\mathbb{R}^{N\times L}$, where $L$ is the dimensionality of the input data. The output layer is represented by the weight matrix $W_{\text{out}}\in\mathbb{R}^{M\times N}$, where $M$ is the number of output neurons. These weight matrices govern the flow of information from the input to the reservoir and from the reservoir to the output layer, ensuring the appropriate transformations and mappings of data within the network. Typically, both $W_{\text{in}}$ and $W_{\text{r}}$ are randomly initialized and remain fixed throughout operation, whereas $W_{\text{out}}$ is optimized via gradient descent or ordinary least-squares methods to obtain the optimal parameters. This strategy substantially reduces the training time and computational resources.
@@
where $\mathbf{u}_{t+1}\in\mathbb{R}^{L}$ is the input vector, $\mathbf{x}_{t}\in\mathbb{R}^{N}$ is the previous reservoir state, $W_{\text{in}}\in\mathbb{R}^{N\times L}$ and $W_{\text{r}}\in\mathbb{R}^{N\times N}$ are the fixed input and reservoir weight matrices, $\mathbf{b}$ is an optional bias, and $\phi(\cdot)$ is a point-wise non-linear activation (e.g., $\tanh$). The leak rate $\alpha$ controls the trade-off between retaining historical information and incorporating new information.
More precisely, the activation of each neuron inside the reservoir at time step $t+1$ depends on (i) the subset of input-layer neurons projecting to it and (ii) the neighbouring reservoir neurons to which it is connected. Denoting the $i$-th reservoir neuron by $x_{i,t}$, we have
@@
Inspection of the update rules in Eqs. (1)–(2) reveals that, much like a conventional recurrent neural network (RNN)[33], the reservoir maintains a compact representation of both historical and instantaneous information within its state $\mathbf{x}_{t}$. This internal memory allows the model to integrate long-range temporal dependencies when predicting future outcomes. The output at time $t$ is produced by the current reservoir state:
@@
In our study, we focus on the Echo State Network (ESN), a widely used framework of reservoir computing. Here, an important index that is closely related to the properties of ESN is the (short-term) memory capacity (MC). The memory capacity quantifies the network’s ability to reconstruct past information from its reservoir at the network’s output by computing correlations [34, 35, 36]. It is defined by the following equation:
@@
Here, $\operatorname{cov}$ represents the covariance between the two time series, and $\operatorname{var}$ denotes the variance. $\mathbf{u}(t-k)$ refers to the input presented $k$ steps before the current input, while $\mathbf{y}_{k}(t)=\mathbf{\tilde{u}}(t-k)$ is its reconstructed output from the network using a linear readout. Thus, memory in ESNs is conceptualized as the network’s capacity to retrieve historical information (for various $k$ values) from the reservoir through linear combinations of internal unit activations. The study [36] shows that the memory capacity of ESNs depends on the properties of the structure and parameters of reservoirs, and may also be influenced by the data properties.
A key finding of our study is that introducing additional rings into the reservoir digraph enhances the orthogonality of its adjacency matrix, a property empirically demonstrated to correlate with improved reservoir performance. The orthogonality of reservoir adjacent matrices is closely related to the memory capability of reservoirs, which influences the prediction of reservoirs [37]. Mathematically, orthogonality is a special case of linear independence, which ensures that the matrix columns are not only linearly independent but also mutually orthogonal. Improving the orthogonality of a reservoir matrix reduces overlap and interference between dataset signals stored at different time steps, enabling the network to better distinguish and retrieve distinct inputs. Moreover, orthogonal matrix columns or rows increase the probability of the matrix being full-rank, which is crucial for maximizing memory capacity. Additionally, enhanced orthogonality improves numerical stability, ensuring that computational errors do not accumulate and allowing the network to store and recall information with greater accuracy. In summary, enhancing the orthogonality of reservoir matrices maximizes short-term memory capacity and improves the network’s ability to distinguish and retrieve different input signals [37]. This provides a theoretical base for our optimization method, which focuses on improving the orthogonality of reservoir matrices.
Hence, we aim to increase the orthogonality of the reservoir adjacent matrices to improve their capabilities. Here we use the following measurement to measure the orthogonality of a matrix:
@@
It is easy to see that $OM(A)$ ranges from $[0,1]$, and the closer to zero, the more vectors in $\{\textbf{a}_{1},\textbf{a}_{2},\cdots,\textbf{a}_{n}\}$ are orthogonal to each other, which implies that the orthogonality of this matrix is greater.
Since a reservoir can be represented by a digraph, the problem of improving the orthogonality of its adjacency matrix can also be regarded from the perspective of the reservoir digraph corresponding to the adjacency matrix. To explain this, we first introduce ring, which is a special kind of digraph and is useful for analyzing the properties of reservoirs. A digraph $G=(V,E)$ is a ring if it is connected (as an undigraph), and the in-degree and out-degree of each vertex are 1. We say a digraph has rings if some of its subgraphs are rings. According to the matrix representation of a ring, modifying a reservoir to increase the number of rings in its digraph is equivalent to making its adjacency matrix possess more fundamental cyclic submatrices like:
@@
or its transpose. For example, in Fig. 4, when changing from (a) to (b), the shape A{A}D{D}B{B}C{C} becomes A{A}D{D}B{B}C{C}, and the corresponding matrix representation changes from $\begin{pmatrix}&1&&1\\ &&1&\\ &&&1\\ &&&\end{pmatrix}$ to $\begin{pmatrix}&1&&\\ &&1&\\ &&&1\\ 1&&&\end{pmatrix}$. Notice that the orthogonality of the adjacency matrix increases from 3 to 4. In fact, for the more general case (Equation 4), since each row and column of this matrix is orthogonal to each other, the more submatrices of this type a matrix has, the greater its orthogonality will be. Combining the previous discussion of the relationship between the orthogonality of the reservoir matrix and the memory capacity of the reservoir, when we make the reservoir matrix (particularly sparse one) have more fundamental cyclic submatrices (which is equivalent to making the digraph of the reservoir have more rings), the greater the memory capacity of the reservoir is, and the more likely it is to exhibit greater predictive ability [37].
Therefore, the core objective of this study is to enhance the orthogonality of the reservoir adjacency matrix. From a graph-theoretic perspective, this goal is equivalent to adding the number of rings in the reservoir digraph—a equivalence that forms the logical starting point for the subsequent methodological design. To achieve this objective, the GLMY homology theory is introduced in this paper as the core analytical tool and operational framework. It is crucial to emphasize that, within the algebraic topological characterization of digraphs, rings are essentially a proper subset of the set of representative cycles contained in the GLMY homology groups. This implies that the representative cycles described by GLMY homology groups include not only rings but also other non-cyclic loop structures.
@@ 4.2 Optimizing reservoir structure by persistent GLMY homology
$\mathbb{H}_{1}=\{H_{1}(G^{\delta})\xrightarrow{i_{\delta,\delta^{\prime}}}H_{1}(G^{\delta^{\prime}})\}_{\delta\leq\delta^{\prime}},$
@@
Then, the one-dimensional GLMY homology of simple digraphs is clear because of the following Theorem [26]. This theorem establishes the theoretical foundation for our operation to reservoirs.
@@
Theorem. Let $G=(V,E)$ be a simple digraph. Let $Z_{1}=\operatorname{Ker}\left(\left.\partial\right|_{\Omega_{1}}\right),B_{1}=\operatorname{Im}\left(\left.\partial\right|_{\Omega_{2}}\right)$ , and let $Q$ denote the space generated by all bigons and boundary triangles and boundary squares in $G$. Then, we have $B_{1}=Q$. Hence the one-dimensional GLMY homology group satisfies that $H_{1}=Z_{1}/Q$.
@@
Consequently, any 1-cycle in a digraph that does not form a bigon, boundary triangle, or boundary square serves as a generator of the one-dimensional GLMY homology. In our study, we employ an algorithm that computes the one-dimensional persistent GLMY homology [26] to identify the minimal generators of $H_{1}(G)$ for a given digraph $G$. These minimal generators are called the minimal representative cycles of the $H_{1}(G)$.
@@
Now we introduce our approach for optimizing reservoir structures using GLMY homology to enhance their predictive capabilities. As we have mentioned above, a key distinguishing feature of reservoirs compared to feedforward neural networks is the presence of rings in their architecture, which have been shown to possibly influence their time series prediction performance. Therefore, we aim to increase the number of rings in a given reservoir to optimize its structure. Specifically, when a reservoir is given, we firstly compute the persistent GLMY homology of its digraph and get the minimal representative cycles of one-dimensional GLMY homology groups. Since ring is a special type of minimal representative cycle, we can modified other representative cycles into rings by changing some edges’ directions in a compatible way (this will be introduced later). In other words, these representative cycles are used to modify the structure of the reservoir to increase the number of rings in the reservoir digraph. In our experiments, the reservoir can get better prediction ability on three tested datasets after adding the rings in the reservoir digraph. The flowchart of our method and experimental procedures is shown in Fig. 1.
@@
From a theoretical perspective, the 1-dimensional persistence diagram (1-PD) plays a crucial role in our methodology. Each point in the 1-PD corresponds to a minimal representative cycle in the digraph. Our first step is to check whether each minimal representative cycle forms a closed loop, where edges are connected end-to-end in a sequential and continuous manner, which we define as a "ring". This verification process is essential because only cycles with a proper ring structure can fully contribute to the reservoir’s predictive capabilities. When a minimal representative cycle fails to meet the ring condition, we employ an edge direction modification process. The primary objective of this process is to transform the non-ring representative cycle into a ring while ensuring that all existing rings within the digraph are preserved. For each individual representative cycle, we begin by attempting to modify its edges to achieve the desired ring structure. However, during the edge direction modification, if there are common edges among the current cycle and existing rings, changing the direction of these shared edges may inadvertently destroy these existing rings (see Fig. 4 for an example). In such cases, where altering the current cycle would be incompatible with the preservation of existing rings, we adopt a conservative approach. Instead of forcing the modification and risking the disruption of the existing rings, we abandon the attempt to change the current cycle and proceed to the next representative cycle in the sequence.
@@
0: A digraph $G=(V,E)$ representing a reservoir.
1: Computing the persistent GLMY homology of $G$ and get the 1-PD.
2: Find the minimal representative cycle for each point in 1-PD. Denote these cycles by $C$.
3:For each cycle in $C$:
4:  If the cycle is already a ring:
5:    Turn to the next cycle.
6:  Else
7:   If the cycle is not a ring but can be modified as a ring without destroying existing rings:
8:     Change some edge directions to make it a ring.
9:   Else
10:    Turn to the next cycle.
11:  EndIf
12:EndFor
12: The optimized digraph $G$ of the reservoir.
Figure 5: Illustration of four datasets, and their Time-delay embedding results in $\mathbb{R}^{3}$ and the one-dimensional persistence diagrams. (a) Result of Mackey-Glass Data. (b) Result of MSO Data. (c) Result of Lorenz System Data. (d) Result of NARMA Data.

### v2

@@ 4 Methodology
In this section, we first introduce some basic concepts and methods of reservoir computing, as well as some theoretical analysis. In particular, we consider optimization by improving the orthogonality of the reservoir digraph matrix, which enhances its memory capacity. To achieve this goal, the GLMY theory will be used. Subsequently, we further introduce the persistent GLMY homology of a digraph and the proposed approach for optimizing the reservoirs.
@@ 4.1 Reservoir computing and its theoretical analysis
Training the weight matrices in deep learning models is typically computationally expensive. In contrast to conventional neural networks, where the weights of all layers are trained, the reservoir computing (RC) architecture only requires the training of the readout layer’s weights. This characteristic significantly reduces the computational complexity, offering a computational efficiency advantage in certain applications. The general structure of the reservoir computing model is illustrated in Figure 3. It primarily consists of three components: the input layer, the reservoir, and the output layer. The input layer is responsible for encoding the incoming data and feeding it into the reservoir. The reservoir consists of a large, fixed, and recurrent network of neurons that project the input data into a high-dimensional dynamic space. The output layer, which is typically trained, maps the high-dimensional representations from the reservoir to the desired output. Generally, the input and output layers are fully connected layers, while the reservoir is represented by a weight matrix $W_{\text{r}}\in\mathbb{R}^{N\times N}$, where $N$ denotes the size of the reservoir. The weight matrix defines the connections between the neurons within the reservoir, determining the structure and dynamics of the recurrent network. Similarly, the input layer $L_{\text{in}}$ and the output layer $L_{\text{output}}$ are also described by their respective weight matrices. Specifically, the input layer is represented by the weight matrix $W_{\text{in}}\in\mathbb{R}^{N\times L}$, where $L$ is the dimensionality of the input data. The output layer is represented by the weight matrix $W_{\text{out}}\in\mathbb{R}^{M\times N}$, where $M$ is the number of output neurons. These weight matrices govern the flow of information from the input to the reservoir and from the reservoir to the output layer, ensuring the appropriate transformations and mappings of data within the network. Typically, both $W_{\text{in}}$ and $W_{\text{r}}$ are randomly initialized and remain fixed throughout operation, whereas $W_{\text{out}}$ is optimized via gradient descent or ordinary least‑squares methods to obtain the optimal parameters. This strategy substantially reduces the training time and computational resources.
@@
where $\mathbf{u}_{t+1}\in\mathbb{R}^{L}$ is the input vector, $\mathbf{x}_{t}\in\mathbb{R}^{N}$ is the previous reservoir state, $W_{\text{in}}\in\mathbb{R}^{N\times L}$ and $W_{\text{r}}\in\mathbb{R}^{N\times N}$ are the fixed input and reservoir weight matrices, $\mathbf{b}$ is an optional bias, and $\phi(\cdot)$ is a point‑wise non‑linear activation (e.g., $\tanh$). The leak rate $\alpha$ controls the trade‑off between retaining historical information and incorporating new information.
More precisely, the activation of each neuron inside the reservoir at time step $t+1$ depends on (i) the subset of input‑layer neurons projecting to it and (ii) the neighbouring reservoir neurons to which it is connected. Denoting the $i$-th reservoir neuron by $x_{i,t}$, we have
@@
Inspection of the update rules in Eqs. (1)–(2) reveals that, much like a conventional recurrent neural network (RNN)[rumelhart1986learning], the reservoir maintains a compact representation of both historical and instantaneous information within its state $\mathbf{x}_{t}$. This internal memory allows the model to integrate long‑range temporal dependencies when predicting future outcomes. The output at time $t$ is produced by the current reservoir state:
@@
In our study, we focus on the Echo State Network (ESN), a widely used framework of reservoir computing. Here, an important index that is closely related to the properties of ESN is the (short-term) memory capacity (MC). The memory capacity quantifies the network’s ability to reconstruct past information from its reservoir at the network’s output by computing correlations [jaeger2001short, farkavs2016computational, baranvcok2014memory]. It is defined by the following equation:
@@
Here, $\operatorname{cov}$ represents the covariance between the two time series, and $\operatorname{var}$ denotes the variance. $\mathbf{u}(t-k)$ refers to the input presented $k$ steps before the current input, while $\mathbf{y}_{k}(t)=\mathbf{\tilde{u}}(t-k)$ is its reconstructed output from the network using a linear readout. Thus, memory in ESNs is conceptualized as the network’s capacity to retrieve historical information (for various $k$ values) from the reservoir through linear combinations of internal unit activations. The study [baranvcok2014memory] shows that the memory capacity of ESNs depends on the properties of the structure and parameters of reservoirs, and may also be influenced by the data properties.
A key finding of our study is that introducing additional rings into the reservoir digraph enhances the orthogonality of its adjacency matrix, a property empirically demonstrated to correlate with improved reservoir performance. The orthogonality of reservoir adjacent matrices is closely related to the memory capability of reservoirs, which influences the prediction of reservoirs [strauss2012design]. Mathematically, orthogonality is a special case of linear independence, which ensures that the matrix columns are not only linearly independent but also mutually orthogonal. Improving the orthogonality of a reservoir matrix reduces overlap and interference between dataset signals stored at different time steps, enabling the network to better distinguish and retrieve distinct inputs. Moreover, orthogonal matrix columns or rows increase the probability of the matrix being full-rank, which is crucial for maximizing memory capacity. Additionally, enhanced orthogonality improves numerical stability, ensuring that computational errors do not accumulate and allowing the network to store and recall information with greater accuracy. In summary, enhancing the orthogonality of reservoir matrices maximizes short-term memory capacity and improves the network’s ability to distinguish and retrieve different input signals [strauss2012design]. This provides a theoretical base for our optimization method, which focuses on improving the orthogonality of reservoir matrices.
Hence, we aim to increase the orthogonality of the reservoir adjacent matrices to improve their capabilities. We use the following measurement to measure the orthogonality of a matrix:
@@
It is easy to see that $OM(A)$ ranges from $[0,1]$, and the closer to zero, the more vectors in $\{\textbf{a}_{1},\textbf{a}_{2},\cdots,\textbf{a}_{n}\}$ are orthogonal to each other, which implies that the orthogonality of this matrix is greater. We clarify that this metric is a normalized adaptation of Mutual Coherence [donoho2001uncertainty] and is closely related to the Frame Total Potential [benedetto2003finite]. In high-dimensional dynamical systems like reservoirs, the degree of linear independence between state dimensions is critical for information processing. Our OM metric quantifies the average “aliasing” between basis vectors. By defining OM, we provide a quantitative tool to track how our GLMY theory-based topological optimization systematically reduces redundancy and approaches the isometry property required for optimal fading memory.
Since a reservoir can be represented by a digraph, the problem of improving the orthogonality of its adjacency matrix can also be regarded from the perspective of the reservoir digraph corresponding to the adjacency matrix. To explain this, we first introduce ring, which is a special kind of digraph and is useful for analyzing the properties of reservoirs. A digraph $G=(V,E)$ is a ring if it is connected (as an undigraph), and the in-degree and out-degree of each vertex are 1. We say a digraph has rings if some of its subgraphs are rings. According to the matrix representation of a ring, modifying a reservoir to increase the number of rings in its digraph is equivalent to making its adjacency matrix possess more fundamental cyclic submatrices, such as:
@@
or its transpose. For example, in Fig. 4, when changing from (a) to (b), the shape A{A}D{D}B{B}C{C} becomes A{A}D{D}B{B}C{C}, and the corresponding matrix representation changes from $\begin{pmatrix}&1&&1\\ &&1&\\ &&&1\\ &&&\end{pmatrix}$ to $\begin{pmatrix}&1&&\\ &&1&\\ &&&1\\ 1&&&\end{pmatrix}$. Notice that the orthogonality of the adjacency matrix increases from 3 to 4. In fact, for the more general case (Equation 4), since each row and column of this matrix is orthogonal to each other, the more submatrices of this type a matrix has, the greater its orthogonality will be. Combining the previous discussion of the relationship between the orthogonality of the reservoir matrix and the memory capacity of the reservoir, when we make the reservoir matrix (particularly sparse one) have more fundamental cyclic submatrices (which is equivalent to making the digraph of the reservoir have more rings), the greater the memory capacity of the reservoir is, and the more likely it is to exhibit greater predictive ability [strauss2012design].
Therefore, the core objective of this study is to enhance the orthogonality of the reservoir adjacency matrix. From a graph-theoretic perspective, this goal is equivalent to adding the number of rings in the reservoir digraph—an equivalence that forms the logical starting point for the subsequent methodological design. To achieve this objective, the GLMY homology theory is introduced in this paper as the core analytical tool and operational framework. It is crucial to emphasize that, within the algebraic topological characterization of digraphs, rings are essentially a proper subset of the set of representative cycles contained in the GLMY homology groups. This implies that the representative cycles described by GLMY homology groups include not only rings but also other non-cyclic loop structures.
@@ 4.2 Optimizing reservoir structure by persistent GLMY homology
$H_{1}=\{H_{1}(G^{\delta})\xrightarrow{i_{\delta,\delta^{\prime}}}H_{1}(G^{\delta^{\prime}})\}_{\delta\leq\delta^{\prime}},$
@@
Then, the one-dimensional GLMY homology of simple digraphs is clear because of the following Theorem [dey2022efficient]. This theorem establishes the theoretical foundation for our operation to reservoirs.
@@
Theorem. Let $G=(V,E)$ be a simple digraph. Let $Z_{1}=\operatorname{Ker}\left(\left.\partial\right|_{\Omega_{1}}\right),B_{1}=\operatorname{Im}\left(\left.\partial\right|_{\Omega_{2}}\right)$ , and let $Q$ denote the space generated by all bigons and boundary triangles and boundary squares in $G$. Then, we have $B_{1}=Q$. Hence, the one-dimensional GLMY homology group satisfies that $H_{1}=Z_{1}/Q$.
@@
Consequently, any 1-cycle in a digraph that does not form a bigon, boundary triangle, or boundary square serves as a generator of the one-dimensional GLMY homology. In our study, we employ an algorithm that computes the one-dimensional persistent GLMY homology [dey2022efficient] to identify the minimal generators of $H_{1}(G)$ for a given digraph $G$. These minimal generators are called the minimal representative cycles[grigor2012homologies, dey2022efficient] of the $H_{1}(G)$.
@@
Now we introduce our approach for optimizing reservoir structures using GLMY homology to enhance their predictive capabilities. As we have mentioned above, a key distinguishing feature of reservoirs compared to feedforward neural networks is the presence of rings in their architecture, which have been shown to possibly influence their time series prediction performance. Therefore, we aim to increase the number of rings in a given reservoir to optimize its structure. Specifically, when a reservoir is given, we first compute the persistent GLMY homology of its digraph and get the minimal representative cycles of one-dimensional GLMY homology groups. Since a ring is a special type of minimal representative cycle, we can modify other representative cycles into rings by changing some edges’ directions in a compatible way (this will be introduced later). In other words, these representative cycles are used to modify the structure of the reservoir to increase the number of rings in the reservoir digraph. In our experiments, the reservoir can get better prediction ability on three tested datasets after adding the rings in the reservoir digraph. The flowchart of our method and experimental procedures is shown in Fig. 1.
@@
From a theoretical perspective, the 1-dimensional persistence diagram (1-PD) plays a crucial role in our methodology. Each point in the 1-PD corresponds to a minimal representative cycle in the digraph. Our first step is to check whether each minimal representative cycle forms a closed loop, where edges are connected end-to-end in a sequential and continuous manner, which we define as a “ring”. This verification process is essential because only cycles with a proper ring structure can fully contribute to the reservoir’s predictive capabilities. When a minimal representative cycle fails to meet the ring condition, we employ an edge direction modification process. The primary objective of this process is to transform the non-ring representative cycle into a ring while ensuring that all existing rings within the digraph are preserved. For each individual representative cycle, we begin by attempting to modify its edges to achieve the desired ring structure. However, during the edge direction modification, if there are common edges among the current cycle and existing rings, changing the direction of these shared edges may inadvertently destroy these existing rings (see Fig. 4 for an example). In such cases, where altering the current cycle would be incompatible with the preservation of existing rings, we adopt a conservative approach. Instead of forcing the modification and risking the disruption of the existing rings, we abandon the attempt to change the current cycle and proceed to the next representative cycle in the sequence.
@@
0: A digraph $G=(V,E)$ with $n$ vertices and $m$ edges representing a reservoir.
1: Computing the minimal representative cycles of the 1D GLMY homology of $G$. Denote these cycles by $C$.
2:For each cycle in $C$:
3:  If the cycle is already a ring:
4:    Turn to the next cycle.
5:  Else
6:   If the cycle is not a ring but can be modified as a ring without destroying existing rings:
7:     Change some edge directions to make it a ring.
8:   Else
9:    Turn to the next cycle.
10:  EndIf
11:EndFor
11: The optimized digraph $G$ of the reservoir.
Now we analyze the complexity of Algorithm 1. According to the theoretical bounds provided by [dey2022efficient], computing 1-dimensional persistent GLMY homology and extracting the minimal representative cycles (the first step of our method) follows the complexity of $O(rm^{\omega-1}+m^{\omega}+nmg^{\omega-1})$, here
$r=\min\{a(G)m,\sum_{(u,v)\in E}(d_{in}(u)+d_{out}(v))\},$
$a(G)=O(n)$ is the so-called arboricity of $G$, $\omega<2.373$ is the exponent for matrix multiplication, and $g=rank(H_{1})$. Upon obtaining the set of $g$ minimal representative cycles, the algorithm performs a targeted structural refinement. This phase is characterized by its linear efficiency relative to the reservoir’s sparsity: For each identified generator, the algorithm verifies its cyclic structure and performs a compatible edge-direction modification. The complexity for each generator is $O(L_{i})$, where $L_{i}$ is the length of the cycle. Summed over all $g$ generators, the total complexity for this stage is $O(\sum_{i=1}^{g}L_{i})$. In practice, often empirically near-linear in $m$, the number of edges in the sparse reservoir. Consequently, under the sparse setting considered in our experiments, the complexity of Algorithm 1 is $O(rm^{\omega-1}+m^{\omega}+nmg^{\omega-1})$, concentrated in the initial topology computation.